RC meeting and Run Plan, Thursday October 27th
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:00pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
- Current Polarization ~ ??%
- Lasers turned off for ~10 minutes for replacement of hose in cooling system. Performed NMR measurement after hall restored to beam permit (~13:30).
- Frequent laser trips yesterday afternoon; polarization dropped from 32.3% at 13:15 to 29.7% 16:11. 6 trips in 30 minute window. Turning off the new lasers (which went in yesterday to replace two that weren't functioning properly) fixed the tripping problem. Plan is to run like this for now (~100W), and revive and restore the two lasers removed yesterday. Update?
- (Polhe1 and Polhe4 need to include Microsoft update exemptions so they do not automatically update and reboot.)
- Beam down for RF recovery, 7:45 to ~13:00. Beam restored at ???
- Several things happened in the hall during this period
==== Hall A Walk-Through==== COMPLETED
- Hall A walk-through training (Jessie and Rachel)
====BCM and Unser==== COMPLETED (Simona performing UNSER calibration - STATUS?)
- (problem description - from yesterday) Keith Cole and Dave Mack: When the beamline loses vacuum and is then pumped back down, BCM & Unser are detuned. Keith Cole could support a retuning of the Unser if there is an RF recovery. Dave Mack: For an asymmetry, the current BCM calibration would be fine. For cross sections.... may want another BCM calibration. During controlled access, Keith Cole's group would need 1 hour to retune the Unser (2 People)
====Moller Q3 polarity==== DECIDED TO DELAY UNTIL MONDAY ACCESS
- Jim Coleman could swap the leads to the Q3 Moller quad. Will take ~45 minutes and require 2 people
====SBS Standalone DAQ==== COMPLETED (STATUS?-Alexandre)
- (If not completed or doesn't work, consider running on a limited basis; day/swing, when experts available, turn off if causing trouble)
- Note, we're currently running with GEMs in BB and SBS using main data acquisition
====SBS GEM gas readout==== COMPLETED (power cycled)
- Jack reset rPi - Binary Gas Analyzer (https://logbooks.jlab.org/entry/4071016)
- "When I went into the hall I found the BGA in a bad state. It was cycling down from a value and then jumping back up and cycling down again. I disconnected the power and then re-connected it. It seems to be stable, now."
- Paul King proposed short test to verify polarization; decided it was not necessary:
"I had shown the asymmetry plot in https://logbooks.jlab.org/entry/4068672 to Kent, Caryn, and Riad during a parity meeting, and they were wondering what could have caused the large asymmetry variation in parity run 9450, and asked if I was certain that the pockels cell was still getting voltage applied properly (i.e. the beam was still polarized). They proposed doing a large charge asymmetry measurement as a means of testing the pockels cell." However, John has checked with Hall B and they are seeing their double-spin asymmetry showing up, so I think it is fine. We probably don't need to try a large charge asymmetry run right now."
- Overnight target motion issues (ladder motion, communication, etc...): Jack J, Arun (came in ~4am), Ellen, and shift crew eventually got it running again (~3 hours lost due to multiple (unrelated??) issues Arun came in). Update from Arun
- Shift summary with links to 7 related log entries: https://logbooks.jlab.org/entry/4070657
- Pump (cooling for laser room) was making a lot of noise; Travis and Jessie replaced hose - MUCH quieter now.
- OPS and RF would like to study the ramp rate in Hall C while observing their ramp rate. This can be done in a dedicated beam studies, in which "parasitic beam" could be delivered to the halls. Jay anticipates the "parasitic beam" would be similar to what we are receiving now for production but likely with a higher trip rate. The test duration would be no more than 4 hours.
- If it cannot be done due to Hall A's ramp rate (1 uA/s), they wonder if the ramp on the polarized helium-3 cell could be raised? To ~1.5 uA/s? Switch to a non-glass cell during this period?
- Estimate of the total fraction of events collected? (from yesterday)
- Accumulated charge corrected for target polarization: 3.9 Coulombs [Where does this number (3.9) come from?]
- Excel spreadsheet created to track the DAQ crashes their sources was setup. [Where - not the one on the Run_Coordinator page]
- Other Subsystem updates?
- Configuration and pass change on Monday (10/31/2022): 
- Swapping of the leads to the power supply of the 3rd Moller quad
- Opportunistic BCM calibration (higher current) before pass change (only if Unser is detuned)?
- Luminosity scans with GEMs (May need to interrupt data taking if we want it for this kinematic setting): Need to call Simona to set beam position on optics target and verify that data looks good, ion chamber rates are OK,....
- "Approved to do SBS GEM lumi scan using C-optics foil when the polarized target is down for say a spin up or some other reason and we cannot do production". - Simona (will/did?) look at optics foil position during beam recovery.
- Prioritize this if NMR shows low polarization after the short laser-off period. Simona is ready to look at the optics target as part of beam recovery (if soon enough)
- Need to decide how much time to spend on this and/or polarization criteria for resuming 3He data taking
- (INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run)
- (Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29)
Current Run Plan
NMR measurement Restore beam (standard post-downtime procedure)
- Check optics target - Simona
- GEM Rate Studies?? GEM expert on call
- GEM HV scan??
When polarization high enough, go to polarized helium-3 production.
Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)
- Keep an eye out for drifting BPM positions
- Log all DAQ crashes in the excel worksheet along with the reason for the crash
- If beam delivery is interrupted for more than 4 hours, perform a checkout of the beam profile and raster follow Simona's procedure: 
- Cross check the rate with what is expected: 
- Cross check with the expected ion chamber readings: 
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
- Shift Openings: Monday (morning) OWL, Tuesday OWL and SWING.